5. The rate constant for the reaction 3A 2B is 0.0447 s What is the half-life of the reaction? (2 points) dbl
Answer #1

By seeing the unit of k, we can say that this is 1st order reaction.

Given:

k = 4.47*10^-2 s-1

use relation between rate constant and half life of 1st order reaction

t1/2 = (ln 2) / k

= 0.693/(k)

= 0.693/(4.47*10^-2)

= 15.5 s

Answer: 15.5 s